If a red kite is 100 ft. off the ground and rises at 8 ft. per second and a blue kite is 1890 ft. off the ground and rises at 5 ft. per second, then how long will it take for the red kite to be higher than the blue kite?

Answer by ankor@dixie-net.com(22740)

If a red kite is 100 ft. off the ground and rises at 8 ft. per second and a
blue kite is 1890 ft. off the ground and rises at 5 ft. per second, then how
long will it take for the red kite to be higher than the blue kite?
:
Let t = rising time of the kites (in seconds)
:
Write a distance equation, (altitude in this case), when the red catches the blue
they will be at the same altitude
:
Red alt = Blue alt
8t + 100 = 5t + 1890
8t - 5t = 1890 - 100
3t = 1790
t = 1790/3
t ~ 597 seconds; 9 min 57 seconds, at the same alt
:
We can say about 10 min for the red kite to be higher than the blue